OpenStudy (anonymous):

The sum of two consecutive odd integers is 244. What is the smaller integer? i know what the answer is, but how do i solve it? i forgot and wanting to know for my over-the-phone test

OpenStudy (anonymous):

i know that the two integers are 121 and 123 but how do i get it correctly?

OpenStudy (anonymous):

x+x+2=244. simplify to get 2x+2=244. then you subtract 2 from both sides, leaving you with 2x=242. then you divide by 2 on both sides, leaving you with x=121. so the smaller integer would be 121. is this what you got?

OpenStudy (yolo_king):

Did you understand ????^^^

OpenStudy (anonymous):

oh i see makes sense

OpenStudy (anonymous):

thanks

OpenStudy (kainui):

If they're consecutive that just means one comes right after the other yeah? So for example, two odd integers that are consecutive would be 3 an 5, another example is 51 and 53 right? Notice that each number is just the same number but it has 2 added to it, see 51+2=53 and 3+2=5. So we can call our unknowns n and (n+2) and so this allows you to solve for n, the lower number and n+2 the higher number after you add them and set this equal to their sum.

OpenStudy (anonymous):

when finding two consecutive odd/even integers you use this (x) + (x + 2) = ? when finding two consecutive integers you use this (x) + (x+1) = ? correct?

OpenStudy (anonymous):

that is how you do it- i dont know what else you want. :P

OpenStudy (anonymous):

finding three odd/even integers is (x) + (x + 2) + (x + 4) = ? and three consecutive numbers is (x) + (x+1) + (x+2) = ? correct?

OpenStudy (anonymous):

no proplem. :)

OpenStudy (anonymous):

ok so what i said is right ^^

OpenStudy (anonymous):

ok just checking thanks all